A sample has a mean of 50 and a standard deviation of 4.0. Find the z-score for each value of x:
a)x = 54

Found 2 solutions by stanbon, robertb:
Answer by stanbon(75887)

A sample has a mean of 50 and a standard deviation of 4.0. Find the z-score for each value of x:
a)x = 54
----
--
z(54) = (54-50)/0.4 = 4/0.4 = 10
---------
Cheers,
Stan H.




Answer by robertb(5830)